# Visatronic: A Multimodal Decoder-Only Model for Speech Synthesis

### v1

@@ Abstract
In this paper, we propose a new task – generating speech from videos of people and their transcripts (VTTS) – to motivate new techniques for multimodal speech generation. This task generalizes the task of generating speech from cropped lip videos, and is also more complicated than the task of generating generic audio clips (e.g., dog barking) from videos and text. Multilingual versions of the task could lead to new techniques for cross-lingual dubbing. We also present a decoder-only multimodal model for this task, which we call Visatronic. This model embeds vision, text and speech directly into the common subspace of a transformer model and uses an autoregressive loss to learn a generative model of discretized mel-spectrograms conditioned on speaker videos and transcripts of their speech. By embedding all modalities into a common subspace, Visatronic can achieve improved results over models that use only text or video as input. Further, it presents a much simpler approach for multimodal speech generation compared to prevailing approaches which rely on lip-detectors and complicated architectures to fuse modalities while producing better results. Since the model is flexible enough to accommodate different ways of ordering inputs as a sequence, we carefully explore different strategies to better understand the best way to propagate information to the generative steps. To facilitate further research on VTTS, we will release (i) our code, (ii) clean transcriptions for the large-scale VoxCeleb2 dataset, and (iii) a standardized evaluation protocol for VTTS incorporating both objective and subjective metrics.
@@ 1 Introduction
Figure 1: Visatronic overview. In addition to existing text-to-speech (left top) and lips-to-speech tasks (right top), we propose a novel multimodal generative task (bottom), video-text-to-speech (VTTS), where the model is conditioned on the video of talking people and corresponding text transcriptions in order to generate speech. Also, we propose a unified multimodal decoder-only architecture, Visatronic, that processes all modalities (video $\mathbf{v}$ (grey), text $\mathbf{t}$ (grey), and speech $\mathbf{s}$ (blue)) in the LM-style transformer model after all modalities are discretized. The model is trained using cross entropy loss $\mathcal{L}_{CE}$ computed on speech discrete values $\mathbf{s}_{t}$ given the mixed multimodal input sequence ${\{\mathbf{v}\}_{t},\{\mathbf{t}\}_{i},\{\mathbf{s}\}_{t}}$. Each input modality is processed in the unified framework, enabling the model to learn interactions between different modalities while learning the temporal alignment.
The research community has made strides in building multimodal models for speech and audio generation. These techniques have been driven by two different types of problems: {NoHyper} **footnotetext: Work done during internship at Apple. generating speech from cropped videos of lips [46], and generating audio (e.g. barking of dogs) from textual descriptions and videos [19]. The former problem simplifies the task of video-conditioned speech generation by using a pretrained model to crop out lips, while the latter deals with generating outputs whose content is very loosely specified and does not need to show as strong correspondence as speech does to a text sequence. In this paper, we propose a new task – generating speech from videos of people and their transcripts (VTTS) – to motivate new techniques for multimodal speech generation. VTTS is more complicated than the above tasks in several different ways. Firstly, the task is defined as an end-to-end task, in that it does not require additional models to detect and crop the lips in the videos. Secondly, the synthesis must satisfy multiple critical criteria: the speech must be clearly intelligible by following the input text, precisely synchronized with the speaker’s movements, and sound natural in terms of prosody and speaking style. In addition, it should leverage facial features that are informative to the task of speech generation, such as emotion and intensity, and also be consistent with other events in the video. We believe VTTS can enable novel applications beyond existing speech generation tasks. For example, multilingual models trained with this approach could be used to perform video dubbing across different languages.
Multimodal generative modeling has made rapid strides recently using auto-regressive transformer models [19, 48, 26] and can be applied to VTTS. These methods piggyback on the observation that transformer-based large language models (LLMs) can learn extremely complicated distributions using next-step prediction. In order to do so, these approaches typically use a vector-quantized variational autoencoder (VQ-VAE) [40] to convert the inputs from the different modalities to sequence of discrete tokens that the language model can consume. Using this recipe, prior work has been able to generate data such as videos, images and speech conditioned on text input [48, 19, 45, 4].
Recently, it has also been shown that a similar autoregressive approach can be used with joint models of text and speech, without tokenization by simply quantizing the mel-spectrogram of speech into discrete, uniformly spaced bins [3]. In this paper, we show that this approach can be generalized and applied to VTTS. We call our model Visatronic. Visatronic embeds each of the modalities – text, vision and speech – into the embedding space of the transformer. Text is input to the model by tokenization followed by embedding lookup. Videos are converted to discrete representations using a VQ-VAE and Visatronic learns to embed them through a special embedding scheme. Speech is quantized and embedded through a similar scheme to the vision inputs.
To evaluate Visatronic’s effectiveness in real-world scenarios, we conduct extensive experiments on the LRS3 [1] dataset following [46], and a more challenging dataset VoxCeleb2 [9], which contains “in-the-wild” videos featuring hundreds of unique speakers with unconstrained vocabulary and diverse acoustic conditions. Compared to LRS3, VoxCeleb2 is 3x larger and contains paired video-speech data without text, has a more diverse and larger pool of speakers and acoustic conditions, and larger background noise. Given these factors, we mainly focus on the VoxCeleb2 dataset in the paper.
To the best of our knowledge, there are no standarized evaluation protocols for VTTS, so we establish a comprehensive evaluation framework that combines both subjective human assessments and objective metrics. Furthermore, we implement and evaluate multiple baseline approaches to provide meaningful comparisons and facilitate future research in this emerging field. Our results demonstrate that Visatronic performs better than prior techniques that use either cropped lips or text as inputs, achieving 12.2% word error rate (WER) on VoxCeleb2 [9] and 4.5% WER on LRS3 [1] datasets. These results also demonstrate that Visatronic generalizes robustly to diverse visual and acoustic conditions not seen during training.
Our contributions are summarized below as follows:
  • We propose a new multimodal generative task, video-text-to-speech (VTTS), to facilitate research in multimodal generation and understand importance of the video conditioning for speech generation.
@@
  • We show the importance of the data processing pipeline to prepare triplets of (video, text, speech) for model training and provide the clean transcriptions for the VoxCeleb2 [9].
  • We successfully trained a unified multimodal decoder-only model for speech generation. We show that conditioning on both video and text improves speech generation over the TTS models across both objective and subjective metrics, e.g. word error rate of a speech recognition model on the generated speech is reduced by more than relative 15%.
@@ 2 Visatronic
Figure 2: Video representation. The input video frame at time $t$ is mapped via encoder of VQ-VAE model to the downsampled spatial representation being in $\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$. Then every spatial location $(h,w)$, with a vector representation in $\mathbb{R}^{D}$, e.g. $(0.8,\dots)$, is mapped to a discrete value, e.g. $3$, using the learned VQ-VAE codebook $\mathbf{C}^{v}$ by finding the closest codebook element via $l_{2}$ similarity. The discrete value at spatial location $(h,w)$ and time $t$ is mapped to a representation $\mathbb{R}^{D^{\prime}}$ via a learnable embedding layer, where $D^{\prime}$ is the transformer decoder dimension. Finally, we perform different aggregations across spatial grid $H^{\prime}\times W^{\prime}$ for embeddings to obtain the final embedding for the frame before inputting it to the transformer decoder: (a) a self-attention mechanism where query $Q$, key $K$, and value $V$ transformations learn spatial relationships within the region, (b) summation or (c) mean pooling across all embeddings to capture aggregate spatial information, (d) max pooling across all embeddings to capture the most salient features, (e) stacking all embeddings followed by a learnable linear projection $\mathbb{R}^{H^{\prime}W^{\prime}D^{\prime}}\to\mathbb{R}^{D^{\prime}}$.
In the rest of the paper, we denote tensors as $\mathbf{x}$ while $\mathbf{x}_{i,...}$ denotes the $(i,...)$-th component of tensor $\mathbf{x}$.
2.1 Video-Text-To-Speech (VTTS)
Video-text-to-speech synthesis (VTTS) can be formulated as follows: given (a) the input video frames of the speaker $\mathbf{x}^{v}\in\mathbb{R}^{T^{v}\times H\times W\times 3}$, where $H$ and $W$ are the spatial video resolution (frame height and width, respectively) and $T^{v}$ is the total number of frames in the video; and (b) text tokens $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$ representing the transcript of speech in the video, where $\mathbf{x}^{t}_{i}\in\textit{Vocabulary}$, $N$ is length of the tokenized transcript, the goal is to generate speech signal $\mathbf{x}^{s}\in\mathbb{R}^{T^{s}}$, where $T^{s}$ is length of speech signal, such that spoken words in speech correspond to the written text $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$, and video and speech are aligned in time.
2.2 Input Representation
@@
To obtain a latent representation of the video input $\mathbf{x}^{v}$, we leverage a pretrained VQ-VAE model [45] which is pre-trained on the general video dataset, Kinetics-600  [5], with codebook $\mathbf{C}^{v}=\{\mathbf{c}_{1}^{v},\mathbf{c}_{2}^{v},\dots,\mathbf{c}_{K^{v}}^{v}\}$ of size $|\mathbf{C}^{v}|=K^{v}$, where $\mathbf{c}_{i}\in\mathbb{R}^{D}$. Using the encoder of the VQ-VAE, we map each input video frame $\mathbf{x}^{v}_{t}\in\mathbb{R}^{H\times W\times 3}$ to a latent vector $\mathbf{y}^{v}_{t}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$ with downsampled resolution $H^{\prime}\times W^{\prime}$. Each spatial element in $\mathbf{y}^{v}_{t}$ is then mapped to the index of its nearest codebook entry $\mathbf{v}_{t,h,w}\in\mathbb{C}^{v}=\{1,\cdots,K^{v}\}$ based on $\ell_{2}$ distance. Thus, every input video frame $\mathbf{x}^{v}_{t}$ is represented as $\mathbf{v}_{t}\in[{\mathbb{C}^{v}}]^{H^{\prime}\times W^{\prime}}$ – a set of discrete values, see Figure 2.
We use the VQ-VAE model to discretize video due to its ability to compress the video representation while preserving both spatial and temporal dynamics crucial for video understanding. Concretely, the pre-trained VQ-VAE model [45] compresses videos with $H\times W=224\times 224$ spatial resolution to $H^{\prime}\times W^{\prime}=16\times 16$ spatial resolution with the codebook dimension $D=3264$ and codebook size $K^{v}=2048$. Although this VQ-VAE model is pre-trained on the general videos, we found it reconstructs speakers videos with sufficient quality to preserve necessary spatial information, see Section D in Appendix.
@@
Figure 3: Speech representation. We follow the speech discretization process from dMel [3]: each continuous mel-filterbank at time $t$ extracted from the raw audio is mapped into a discrete values using a codebook of evenly spaced values in the range $[m,M]$ by taking the closest codebook value, where $m$ and $M$ are the minimum and maximum values of log mel-filterbanks computed across the dataset. Afterwards, each discretized log mel-filterbank at time $t$ is mapped through a learnable embedding layer, all representations for log mel-filterbanks at time $t$ are stacked together and resulting vector is projected by a learnable linear layer to the model dimention $D^{\prime}$. The example illustrates this process by converting log mel-filterbank values $(5.1,2.8,-0.4,\dots)$ into bin indices $(10,8,2,\dots)$, which are then embedded for the Visatronic model processing. Note, that all discretized log mel-filterbanks at time $t$ are predicted in parallel and independently by a decoder-only model.
Following quantization, every discrete value is mapped via a learnable embedding layer $\mathbf{E}^{v}(\cdot):\mathbb{C}^{v}\to\mathbb{R}^{D^{\prime}}$ to $\mathbf{e}^{v}_{t,h,w}$. The representation for the whole frame after embedding is $\mathbf{e}_{t}^{v}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D^{\prime}}$, where $D^{\prime}$ is the Visatronic transformer input dimension. Subsequently, we explore various methods for aggregating the spatial dimensions of this representation prior to inputting to the transformer decoder:
Max pooling: $\mathbf{z}_{t}^{v}=\max_{(h,w)}\mathbf{e}^{v}_{t,h,w}$;
Stacking: stack embeddings and then project it via a learnable linear layer $\mathbf{L}^{v}(\cdot):\mathbb{R}^{H^{\prime}W^{\prime}D^{\prime}}\to\mathbb{R}^{D^{\prime}}$,
$\mathbf{z}_{t}^{v}=\mathbf{L}^{v}([\mathbf{e}^{v}_{t,1,1},\mathbf{e}^{v}_{t,1,2},\dots,\mathbf{e}^{v}_{t,1,W^{\prime}},\mathbf{e}^{v}_{t,2,1},\dots,\mathbf{e}^{v}_{t,H^{\prime},W^{\prime}}])$.
As we show later, this multi-faceted approach enables effective capture of both local and global video characteristics in Visatronic.
@@
Text Representation    For text processing, we employ a character-level tokenizer that maps the input text $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$ to a sequence of discrete tokens $\mathbf{t}_{j}\in\mathbb{C}^{t}=\{1,2,\dots,K^{t}\}$ with $|\mathbb{C}^{t}|=K^{t}$, followed by a learnable embedding layer $\mathbf{E}^{t}(\cdot):\mathbb{C}^{t}\to\mathbb{R}^{D^{\prime}}$. Character-level tokenization reduces vocabulary size $K^{t}$ and improves generalization by capturing fine-grained linguistic features.
@@
Speech Representation    We utilize dMel [3], a simple yet effective discretization approach for speech processing; see Figure 3 for an overview. Given an input speech signal $\mathbf{x}^{s}$, we first compute continuous log mel-filterbanks $\mathbf{y}_{t}^{s}\in\mathbb{R}^{F}$ for a frame at time $t$, where $F$ is the number of log mel-filterbanks. Then, we map every log mel-filterbank $\mathbf{y}_{t,f}^{s}\in\mathbb{R}$ to a discrete value $\mathbf{s}_{t,f}\in\mathbb{C}^{s}=\{1,2,\dots 2^{K^{s}}\}$ using a codebook $\mathbf{C}^{s}=\{\mathbf{c}^{s}_{1},\mathbf{c}^{s}_{2},\dots,\mathbf{c}^{s}_{2^{K^{s}}}\}$: $\mathbf{c}^{s}_{i}\in\mathbb{R}$ are evenly spaced values in the range $[m,M]$, where $m$ and $M$ are the minimum and maximum values of mel-filterbanks computed across the dataset. To discretize, we take the closest codebook value, i.e., $\mathbf{s}_{t,f}=\text{argmin}_{i\in\mathbb{C}^{s}}|\mathbf{y}_{t,f}^{s}-\mathbf{c}^{s}_{i}|$.
After each speech frame is discretized, every discrete value is mapped via a learnable embedding layer $\mathbf{E}^{s}(\cdot):\mathbb{C}^{s}\to\mathbb{R}^{d^{\prime}}$ to a representation $\mathbf{e}^{s}_{t,f}$. The representation for the whole frame is given by $\mathbf{e}_{t}^{s}\in\mathbb{R}^{F\times d^{\prime}}$, where $d^{\prime}$ is the intermediate dimension. Subsequently, we stack these embeddings and project the resulting vector to a final embedding $\mathbf{z}^{s}_{t}\in\mathbb{R}^{D^{\prime}}$ via a learnable linear layer $\mathbf{L}^{s}(\cdot):\mathbb{R}^{Fd^{\prime}}\to\mathbb{R}^{D^{\prime}}$: $\mathbf{z}_{t}^{s}=\mathbf{L}^{s}([\mathbf{e}^{s}_{t,1},\mathbf{e}^{s}_{t,2},\dots,\mathbf{e}^{s}_{t,F}])$.
This training-free discretization enables effective processing of speech signals in our framework. Following [3] we use $K^{s}=4$ bits with $|\mathbb{C}^{s}|=16$, $F=80$ log mel-filterbank channels and $d^{\prime}=24$.
Speech Inversion    To reconstruct the speech signal $\mathbf{x}^{s}$ from the speech discrete values $\mathbf{s}_{t,f}$ predicted by the multimodal transformer decoder (Section 2.3), we follow [3]: first, we transform the indices back to the log mel-filterbanks via the codebook $\mathbf{C}^{s}$: $\mathbf{\hat{y}}^{s}_{t,f}=\mathbf{c}^{s}_{\mathbf{s}_{t,f}}$. Subsequently, we apply a vocoder [44] to transform reconstructed log mel-filterbanks $\mathbf{\hat{y}}^{s}_{t,f}$ back into the time domain signal $\mathbf{x}^{s}$. The vocoder is trained independently and is not part of the Visatronic transformer decoder-based model.

### v2

@@ Abstract
The rapid progress of foundation models and large language models (LLMs) has fueled significantly improvement in the capabilities of machine learning systems that benefit from mutlimodal input data. However, existing multimodal models are predominantly built on top of pre-trained LLMs, which can limit accurate modeling of temporal dependencies across other modalities and thus limit the model’s ability to jointly process and leverage multimodal inputs. To specifically investigate the alignment of text, video, and speech modalities in LLM-style (decoder-only) models, we consider a simplified multimodal generation task, Video-Text to Speech (VTTS): speech generation conditioned on both its corresponding text and video of talking people. The ultimate goal is to generate speech that not only follows the text but also aligns temporally with the video and is consistent with the facial expressions. In this paper, we first introduce Visatronic, a unified multimodal decoder-only transformer model that adopts an LLM-style architecture to embed visual, textual, and speech inputs into a shared subspace, treating all modalities as temporally aligned token streams. Next, we carefully explore different token mixing strategies to understand the best way to propagate information from the steps where video and text conditioning is input to the steps where the audio is generated. We extensively evaluate Visatronic on the challenging VoxCeleb2 dataset and demonstrate zero-shot generalization to LRS3, where Visatronic, trained on VoxCeleb2, achieves a 4.5% WER, outperforming prior SOTA methods trained only on LRS3, which report a 21.4% WER. This highlights significant gains across objective metrics, such as word error rate and phoneme-level synchronization, and subjective assessments of naturalness and expressiveness. Additionally, we propose a new objective metric, TimeSync, specifically designed to measure phoneme-level temporal alignment between generated and reference speech, further ensuring synchronization quality. The demo is available at https://apple.github.io/visatronic-demo/.
@@ 1 Introduction
Figure 1: Visatronic overview. In addition to existing text to speech (leftmost) and lips to speech tasks (middle), we address multimodal generative task (rightmost), video-text to speech (VTTS), where the model is conditioned on the video of talking people and corresponding text transcriptions in order to generate speech. Visatronic is a unified decoder-only transformer that processes video $\mathbf{v}$ (grey), text $\mathbf{t}$ (grey), and speech $\mathbf{s}$ (blue) as discrete tokens in a shared sequence. The model is trained with cross-entropy loss $\mathcal{L}_{CE}$ to predict speech tokens, learning cross-modal interactions and temporal alignment.
Recent advances in multimodal generation, e.g. [25, 57, 33, 52, 46], have demonstrated the potential of LLM-style architectures by encoding different modalities and integrating them within a unified decoder-only model. However, as they are built on top of pre-trained LLMs, these models face challenges in accurately modeling temporal dependencies across modalities. For example, when combining text, video, and speech together they often under-performing unimodal models or task specialist models [42, 25, 52]. Unimodal models significantly improved their generation quality in past years, e.g. for videos, images and speech conditioned on text input [57, 25, 54, 5], thereby widening the performance gap with multimodal approaches.
In this paper we aim to study how different modalities can be made to convey more fine-grained information to each other (e.g. through proper temporal alignment) in multimodal models within the LLM-style decoder-only architecture. We focus our study on speech generation conditioned on text and video (video-text to speech, VTTS). This task is simpler than some recent multimodal generative models which generate multiple modalities but more complicated in that it requires fine-grained consistency between the different modalities to solve the task appropriately.
We propose Visatronic, a unified multimodal decoder-only transformer model for VTTS. Visatronic processes video, text, and speech as a single token sequence in a shared embedding space. Text is embedded via standard tokenization, video frames are discretized using a pretrained VQ-VAE model, and speech is quantized into uniformly spaced bins. All inputs are aligned temporally and autoregressively decoded to predict speech tokens. This simple and scalable architecture allows effective joint modeling of video, text, and speech without specialized encoders or alignment modules. We evaluate Visatronic on the LRS3 [1] and VoxCeleb2 [10] datasets, with the latter presenting challenging conditions due to its noisy, multilingual, and diverse speaker pool and containing video-speech paired data with no text. Visatronic achieves 12.2% WER on VoxCeleb2 and 4.5% WER on LRS3 in a zero-shot setting, significantly outperforming prior baselines such as text-only TTS and lip-only models. Additionally, we introduce TimeSync, a new phoneme-level synchronization metric, to assess alignment quality. To the best of our knowledge, this is the first work to formalize and benchmark the VTTS (Video-Text to Speech) task using a single decoder-only transformer model that autoregressively generates speech from multimodal token sequences, without reliance on reference audio, prosody supervision, or handcrafted alignment. Our contributions are summarized below:
@@
  • We develop a data processing pipeline to prepare triplets of (video, text, speech) for model training and provide the clean transcriptions for the VoxCeleb2 [10].
  • We introduce a unified multimodal decoder-only transformer architecture that handles all modalities through a single token space, demonstrating that joint conditioning on both video and text substantially improves speech generation over the TTS baselines, reducing WER by more than 15% relatively.
@@ 2 Visatronic
Figure 2: Video representation. Each video frame at time $t$ is encoded via a VQ-VAE [54] into a downsampled spatial grid in $\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$. Each vector at location $(h,w)$ is quantized to a discrete token using the learned codebook $\mathbf{C}^{v}$ via $l_{2}$ similarity. These discrete tokens are embedded into $\mathbb{R}^{D^{\prime}}$ and aggregated across the spatial grid to produce the final frame-level embedding input to the transformer. See Section 2.2 for details.
2.1 Video-Text To Speech (VTTS)
Video-text to speech synthesis can be formulated as follows: given (a) the input video frames of the speaker $\mathbf{x}^{v}\in\mathbb{R}^{T^{v}\times H\times W\times 3}$, where $H$ and $W$ are the spatial video resolution (frame height and width, respectively) and $T^{v}$ is the total number of frames in the video; and (b) text tokens $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$ representing the transcript of speech in the video, where $\mathbf{x}^{t}_{i}\in\textit{Vocabulary}$, $N$ is length of the tokenized transcript, the goal is to generate speech signal $\mathbf{x}^{s}\in\mathbb{R}^{T^{s}}$, where $T^{s}$ is length of speech signal, such that spoken words in speech correspond to the written text $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$, and video and speech are aligned in time.
2.2 Discrete Input Representation Design
@@
To obtain a latent representation of the video input [^1] In the rest of the paper, we denote tensors as $\mathbf{x}$ while $\mathbf{x}_{i,...}$ denotes the $(i,...)$-th component of tensor $\mathbf{x}$. $\mathbf{x}^{v}$, we use a pretrained VQ-VAE model [54], trained on the Kinetics-600 dataset [6]. Each frame $\mathbf{x}^{v}_{t}\in\mathbb{R}^{H\times W\times 3}$ is encoded to a spatially downsampled latent representation $\mathbf{y}^{v}_{t}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$. Each spatial element in $\mathbf{y}^{v}_{t}$ is then quantized to the nearest entry in the learned codebook $\mathbf{C}^{v}=\{\mathbf{c}_{1}^{v},\dots,\mathbf{c}_{K^{v}}^{v}\}$ using $\ell_{2}$ distance, resulting in a discrete token grid $\mathbf{v}_{t}\in[\mathbb{C}^{v}]^{H^{\prime}\times W^{\prime}}$. This quantization process is illustrated in Figure 2. Although trained on general video data, the VQ-VAE compresses each $224\times 224$ frame into a $16\times 16$ grid using a codebook of size $K^{v}=2048$ and embedding dimension $D=3264$. These discrete tokens are then mapped to continuous embeddings $\mathbf{e}^{v}_{t,h,w}\in\mathbb{R}^{D^{\prime}}$ via a learnable embedding layer $\mathbf{E}^{v}:\mathbb{C}^{v}\rightarrow\mathbb{R}^{D^{\prime}}$, resulting in an embedding map $\mathbf{e}_{t}^{v}\in\mathbb{R}^{16\times 16\times D^{\prime}}$. To convert this grid into a frame-level representation $\mathbf{z}_{t}^{v}\in\mathbb{R}^{D^{\prime}}$ suitable for input to the transformer decoder, we apply a spatial aggregation function over the $16\times 16$ grid: $\mathbf{z}_{t}^{v}=\sum_{h=1}^{16}\sum_{w=1}^{16}\mathbf{e}^{v}_{t,h,w}$. This final vector $\mathbf{z}_{t}^{v}$ serves as a compact, information-rich embedding of the frame that preserves spatial and speaker-specific features while aligning to the decoder’s input dimension. We also explore alternative aggregation strategies, such as attention-based pooling and learned convolutions, detailed in Appendix.
@@
Figure 3: Speech representation. We follow the speech discretization process from dMel [4]: each continuous mel-filterbank at time $t$ extracted from the raw audio is mapped into a discrete values using a codebook of evenly spaced values. Afterwards, each discretized log mel-filterbank at time $t$ is mapped through a learnable embedding layer, all representations for log mel-filterbanks at time $t$ are stacked together and resulting vector is projected by a learnable linear layer to the model dimension $D^{\prime}$. All discretized log mel-filterbanks at time $t$ are predicted in parallel and independently.
Text Representation
@@
For text processing, we employ a character-level tokenizer that maps the input text $\{\mathbf{x}^{t}_{i}\}_{1}^{N}$ to a sequence of discrete tokens $\mathbf{t}_{j}\in\mathbb{C}^{t}=\{1,2,\dots,K^{t}\}$ with $|\mathbb{C}^{t}|=K^{t}$, followed by a learnable embedding layer $\mathbf{E}^{t}(\cdot):\mathbb{C}^{t}\to\mathbb{R}^{D^{\prime}}$. Character-level tokenization reduces vocabulary size $K^{t}$ and improves generalization by capturing fine-grained linguistic features.
@@
Speech Representation    We utilize dMel [4], a simple yet effective discretization approach for speech processing; see Figure 3 for an overview. Given an input speech signal $\mathbf{x}^{s}$, we first compute continuous log mel-filterbanks $\mathbf{y}_{t}^{s}\in\mathbb{R}^{F}$ for a frame at time $t$, where $F$ is the number of log mel-filterbanks. Then, we map every log mel-filterbank $\mathbf{y}_{t,f}^{s}\in\mathbb{R}$ to a discrete value $\mathbf{s}_{t,f}\in\mathbb{C}^{s}=\{1,2,\dots 2^{K^{s}}\}$ using a codebook $\mathbf{C}^{s}=\{\mathbf{c}^{s}_{1},\mathbf{c}^{s}_{2},\dots,\mathbf{c}^{s}_{2^{K^{s}}}\}$: $\mathbf{c}^{s}_{i}\in\mathbb{R}$ are evenly spaced values in the range $[m,M]$, where $m$ and $M$ are the minimum and maximum values of mel-filterbanks computed across the dataset. To discretize, we take the closest codebook value, i.e, $\mathbf{s}_{t,f}=\text{argmin}_{i\in\mathbb{C}^{s}}|\mathbf{y}_{t,f}^{s}-\mathbf{c}^{s}_{i}|$. After each speech frame is discretized, every discrete value is mapped via a learnable embedding layer $\mathbf{E}^{s}(\cdot):\mathbb{C}^{s}\to\mathbb{R}^{d^{\prime}}$ to a representation $\mathbf{e}^{s}_{t,f}$. The representation for the whole frame is given by $\mathbf{e}_{t}^{s}\in\mathbb{R}^{F\times d^{\prime}}$, where $d^{\prime}$ is the intermediate dimension. Subsequently, we stack these embeddings and project the resulting vector to a final embedding $\mathbf{z}^{s}_{t}\in\mathbb{R}^{D^{\prime}}$ via a learnable linear layer $\mathbf{L}^{s}(\cdot):\mathbb{R}^{Fd^{\prime}}\to\mathbb{R}^{D^{\prime}}$: $\mathbf{z}_{t}^{s}=\mathbf{L}^{s}([\mathbf{e}^{s}_{t,1},\mathbf{e}^{s}_{t,2},\dots,\mathbf{e}^{s}_{t,F}])$. This training-free discretization enables effective processing of speech signals in our framework. Following [4] we use $K^{s}=4$ bits with $|\mathbb{C}^{s}|=16$, $F=80$ log mel-filterbank channels and $d^{\prime}=24$.
Speech Inversion    To reconstruct the speech signal $\mathbf{x}^{s}$ from the speech discrete values $\mathbf{s}_{t,f}$ predicted by the multimodal transformer decoder (Section 2.3), we follow [4]: first, we transform the indices back to the log mel-filterbanks via the codebook $\mathbf{C}^{s}$: $\mathbf{\hat{y}}^{s}_{t,f}=\mathbf{c}^{s}_{\mathbf{s}_{t,f}}$. Subsequently, we apply a vocoder [53] to transform reconstructed log mel-filterbanks $\mathbf{\hat{y}}^{s}_{t,f}$ back into the time domain signal $\mathbf{x}^{s}$. The vocoder is trained independently and is not part of the Visatronic model.